OpenStudy (anonymous):

find the x-andy-intercepts 7x-8y=28

OpenStudy (anonymous):

y-intercept: found when x is zero. x-intercept: found when y is zero. 7*0 - 8y =28 -8y =28 y = - 3.5 7x - 8 * 0 = 28 7x = 28 x= 4

OpenStudy (anonymous):

thanks i was working to see if i got what you got and i did thanks

OpenStudy (anonymous):

You're welcome.